OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x3 – 35x2 + 26x – 5 ?

OpenStudy (anonymous):

The zeros are the values of x that make the function equal to zero (also called a root). Here because the polynomial has a degree of 3, we will have three zeros. First, lets set the function equal to zero. \[6x^3 - 35x^2 +26x - 5 = 0\]Let's move the five over, factor out an x, and obtain the following\[x \left( 6x^2 - 35x + 26 \right) = 5\]We can see that one of the zeros must be equal to 5. Now let's solve the remaining second-degree polynomial to obtain the other two zeros. We can use the quadratic equation\[x = {-b \pm \sqrt{b^2 - 4ac} \over 2a }\]

OpenStudy (anonymous):

is this the answer ? 5,1/3 , –1/2

OpenStudy (anonymous):

Check the signs of your roots.

OpenStudy (anonymous):

\[6 x^3-35 x^2+26 x-5=(x-5) (2 x-1) (3 x-1) \]

OpenStudy (anonymous):

5, 1/3, 1/2